what are the values of w,x,y, and z in these equations?
1. w + x + y + z = 2
2. 2w - x - y +2z = 7
3. 2w + 3x + 2y - 2z = -2
4. 3w - 2x - y -3z = -2

Answer by ankor@dixie-net.com(12700)

:
1. w + x + y + z = 2
2. 2w - x - y +2z = 7
3. 2w + 3x + 2y - 2z = -2
4. 3w - 2x - y -3z = -2
:
This has to be solved as a matrix, a tedious task unless you use the matrix feature on your Ti83, or similar calc which I am doing here.
:
a 4 X 5 matrix:
:
+1 +1 +1 +1 +2
+2 -1 -1 +2 +7
+2 +3 +2 -2 -2
+3 -2 -1 -3 -2
Solution:
w = 1.2; x = 1.2; y = -2.2; z = 1.8
:
Check solutions in the 1st equation:
1.2 + 1.2 - 2.2 + 1.8 = +2
:
Check solutions in the last equation:
3(1.2) - 2(1.2) - (-2.2) - 3(1.8) =
3.6 - 2.4 + 2.2 - 5.4 = -2
: